Modern Narrative Orality: Urban Legends, News, and Jokes

Matthew Gray

May 15, 1997

Introduction

Typically, when “media” is discussed, the focus is on technological media, such as the radio, television, films, and recently, software and the World Wide Web. Additionally, “media” is often equated with “mass media” because less well distributed media are less able to have substantial cultural impact. Television, for example, is especially well equipped to very widely deliver a message that has a substantial cultural impact.

Before the existence of these technological innovations, the number of media to choose from was far more restricted. Printed media and oral media still existed, however. In modern culture, it has been argued, because of the economic underpinnings of mass media, cultural narratives have become almost entirely corporate property. While clearly a large amount of modern cultural narrative is generated and owned by large corporations, particularly in the realm of television and films, modern culture is not without narratives that are communally owned. These narratives are primarily transmitted via oral media.

1 A Definition of Oral Media

Oral media here is taken to be somewhat different from how it would typically defined. It is modified to include all communication that is typically done individual-to-individual, or individual-to-small-group. This is a very inclusive definition, and includes phone calls, email, postal mail, newsletters, community announcements, and any other communication where the people involved know each other personally at some level. This definition is not meant to preclude oral media being a “mass media”, it is merely not centralized mass media.

Given this broad definition of oral media, it is pertinent to ask what the significant forms oral media takes in modern culture. The examples above of phone calls, email, postal mail, newsletters, community announcements sound be expanded to include community bulletin boards, water cooler chats at work, other conversation, small scale performance, teaching, preaching, and most professional relationships.

Content of Oral Media

Oral media carries a wide range of content, ranging from gossip to education to logistical information. The three types of content discussed here are “Urban Legends”, News, and Jokes.

Urban Legends (ULs) are perhaps the most generic of the three categories, as narratives which spread orally. Typically, they are false, but not always. They have many of the traditional features of legends: moral lessons, humor, fear, variant.

News has become utterly ubiquitous in society, to the point of there being multiple television stations exclusively dedicated to news content. This fact has obscured the fact that much of news is still transmitted orally. In this context, news is meant to include the speculation and opinions that surround discussion of the actual news content.

Jokes are the final form of oral content discussed here. Jokes have always been a primarily orally transmitted content, even in the literal sense. Television comedy has become one source of transmission for jokes, but typically that is not actually jokes and riddles, but situational comedy. While jokes continue to be transmitted largely verbally, email has become a popular mechanism for the distribution of jokes, in the form of “funny posts” that make their rounds on the Internet.

Each of these forms are interrelated. Urban Legends are often originated as news or are based on news. In many cases, urban legends, either through people mistakenly believing them or through the sheer magnitude of their propagation become news themselves. Urban legends often take the form of jokes as well, and it has been hypothesized that some urban legends got their beginnings as a misinterpreted joke. News, while infrequently making use of jokes in the sense discussed here, is often the topic of jokes and a repeated topic of humor.

Urban Legends

The aforementioned narratives are commonly referred to as “Urban Legends”. Urban Legends are typically transmitted via oral media.

Urban Legends typically play to some strong emotion and often, but not always, have some sort of moral message. Usually, for successful transmission, they appeal to some sense of “public good”.

Craig, Mrs. Fields, and the Ankle Slashers

Briefly, presented here are capsule descriptions of three fairly widespread Urban Legends which will be discussed in more detail.

Craig Shergold. Craig Shergold is a poor little British boy dying of cancer and through the “Make-A-Wish” foundation has wished to get in the Guiness Book of World Records for receiving the most Get Well cards. Please send him cards and make a little boy happy.^{1}

Mrs. Fields. A friend of a friend of mine was buying a bunch of Mrs. Fields cookies, and asked the manager if it would be possible to get the recipe. He said no. My friend then asked if she could buy the recipe and was told “Sure, for two fifty”. She had him add it to her bill and took her cookies and recipe home. Upon receiving her credit card bill, she found it was for over $260 dollars, as she had been charged $250 for the recipe. In retaliation for being duped, she is distributing the recipe everywhere, and here it is....^{2}

Ankle Slashers. In mall parking lots across America, a new type of crime is occurring. Young hoodlums will hide underneath cars in parking lots, and when the shoppers return, the hoodlums will slash their ankles from beneath the car, and steal their purchases and sometimes their car.^{3}

News

Three recent news events will be focused on, and are summarized here.

The OJ Simpson Trial is perhaps one of the most extreme cases of a news feeding frenzy in recent years. OJ Simpson, football player, actor, and hero to many people in the 30-40 age group, was accused of murdering his ex-wife and her lover. He was found not guilty in the criminal trial, despite popular belief that he was, in fact, guilty.

1Entirely true, but incomplete. See Impact of Oral Media section for details

2Entirely false

3Entirely false

The TWA Flight 800 Disaster was a recent air disaster when flight 800 exploded south of Long Island, New York. The cause for the explosion is undetermined and the topic of extensive speculation.

The Sheep Cloning was a relatively straightforward bit of science news that succeeded in hitting an emotional chord, sparking extended, ongoing debate on the ethical implications.

Jokes

While the topic of jokes is clearly vast, this paper will focus on two sorts of riddles and one category of anecdotal humor. They are briefly described here.

OJ Simpson Jokes proliferated extensively during and after the abovementioned trial. They focused on the presumption he was guilty contrasted with his fame.

Blond Jokes have manifested themselves in many forms. They are essentially generic “stupid people” jokes that are often applied to ethnic groups instead of blonds.

“True” Stories are closely related to urban legends. These are comical anecdotes which are claimed to be true. Sometimes it is apparent that it is not true at the punch line, while other times, the truth, or alleged truth adds to the humor.

Impact of Oral Media

By the fact that each of these content types are transmitted orally, they are subject to error and variation more than narratives that are centrally distributed. Variations on each of the above stories exist in substantial variety. For the Craig Shergold story, the boy’s age changes, the boy’s name changes, and whether he wants get well, post, or business cards changes. For news stories, details are added, removed, misinterpreted and exaggerated. In the case of the Flight 800 disaster, speculation about the cause rapidly got restated as fact in other contexts, and details were embellished often when they weren’t available. Jokes constantly change, particularly Blond jokes, often in target depending on what group is doing the disparaging and which group is being targeted.

Urban legends have the additional problem that because of the nature of their transmission, these narratives, all of which typically are reported as true, are very difficult to correct when errors are realized. In the latter two stories, these errors are relatively harmless and merely add color to the story, and at worst, cause misplaced emotions toward entities in the stories. In the case of Craig Shergold however, the inability to correct errors has caused substantial problems.

The true story of Craig Shergold is more or less as is related above, but has the added features that the wish was made in 1989, Craig made it into the Guiness Book in 1990, he received so much attention through his wish, an American philanthropist payed for him to receive medical treatment in the US, and was cured of his cancer in 1991. The mail, however, keeps coming, placing substantial burden on the local Post Offices and the Makea-Wish Foundation. It is estimated Craig has been sent in excess of 100 million cards, though official counting ceased at 60 million.

Often, corrections succeed in rapidly following the legends, lest someone believe them true and take rash or inappropriate actions, but the nature of the media makes it very difficult to do so under some circumstances.

News stories have the added issue of being very solidly based in fact, yet due to their typically one directional way of being started, speculation and misinterpretation run wild.

Particularly in the case of sheep cloning case, where much of the relevant content of the story is outside the knowledge and experience of the average person, speculation and misinterpretation were rampant. Widespread misunderstanding of what “cloning” actually is resulted in popular discussion of ethical issues more disconnected from reality than many realized. Ethical statements propagated that would imply that twins were “unethical”, because they are in fact much like clones would be. Other more informed ethical issues were raised, based on the speculation and claims that human cloning may now be substantially more feasible.

Jokes, unlike the other two forms, are hindered by inadvertent changes or misinterpretations. Legends and news tend to continue to propagate, while jokes often fail to remain funny if they are modified at all. This causes jokes to remain relatively constant, or when they change, to change either drastically or in altogether incidental details.

Cultural Motivations

Of the urban legends, each has its own reasons for propagation. Spread of the Craig Shergold story seems substantially spurred by a combination of compassion and a sense of being able to relatively easily help in what in many ways seems like a helpless situation. As evidenced by the extent to which this legend has spread, it clearly strikes a particular chord within society: A desire to help, especially when it is easy.

The Mrs. Fields story, on the other hand, appeals to a sense of justice. Specifically, it represents a sense of justice that succeeds in vindicating the individual in the face of corporate giants. Once this justice has been served, the story propagates by the desire to share this triumph, and in this case booty, in the form of the recipe.

The ankle slashers story is somewhat less widespread than the others, but is typical of a third brand of fear-motivated legends. Other often recognized legends in this same category include the “laptop thieves at airport security” story, “involuntary organ donation” story, and the “death by jeans shrinking in rain” story. The motivation for distribution of these stories tends to be in the name of enhanced public awareness. In fact most of these stories are entirely false, but the perception of needing to be warned of these threats to our person, particularly in gruesome or otherwise extreme ways, encourages propagation and highlights the specific nature of some the fears society has.

The cultural motivations for the oral transmission of news has long been because it affects people’s lives, directly. For local news this is clearly the case, but it is less so for national or global news. The Cold War, however, changed that. World events became of deep individual importance. A wrong turn in international relations could result in death for the majority of humanity. This motivation extended to other forms of non-local news to the point where news can be of interest even when it has a very tenuous impact on the reader’s life.

Furthermore, knowledge of current events has become something of a social badge of intelligence, or at least “togetherness”. The value of this knowledge as a sheer social commodity has become substantial in making news a truly oral media. While it is natural to discuss life threatening news in oral media, when it becomes a symbol of social stature to discuss so, people are doubly motivated.

The social motivation for the oral transmission of jokes is similar to this last motivation for news. Conveying humor, and hence displaying wit, is almost universally considered a social virtue and strength. Occasionally other motivations apply, including the desire to attack a individual or group in a less direct way than they might otherwise, or the desire to address an issue that cannot be addressed directly.

Additionally, as in the case of OJ Simpson jokes, jokes can be used to convey the dual social qualities of being well versed in current events and being witty.

Technological Impact

The urban legends are relatively modern, and hence seemingly less subject to technological impact, they have in fact come at a time where computer mediated oral communication is becoming widespread.

The Craig Shergold legend, because its basis in fact is well documented, is a case that can be tracked well. The story began its disemmination in 1989, and the actual request was no longer being sent in 1990, and Craig was cured in 1991, all before the recent explosive growth of computer mediated communication. As people began to populate the Internet, substantially driven by the web, new types of oral media became available for distribution of these stories.

The time line of these events identifies two important points. First, the legend of Craig Shergold spread quite well between 1989 and 1991 via oral media not dominated by email or other computer mediated communication. In fact, in this time over 10 million cards were received.

Second, the widespread use of email spread the story even further. The Make-A-Wish foundation has an 800 number to handle questions about Craig Shergold and noticed a roughly four-fold increase in call volume in 1995, as the Internet craze had recently reached true popular acceptance.^{4}

News has been strikingly subject to the impact of technology. Obvious is the technological improvements in mass media presentations of news, but technology has also impacted its oral transmission. With the advent of trivial, instantaneous long distance communication, orality has become again a way to transmit news often before it can be received from the mass media. Major news typically originates from a mass media source, but is then transmitted orally to many people via those readers/viewers.

With the advent of Internet technologies, news has become discussed orally through newsgroups and chat rooms in ways that were not previously possible, such as live conversations between people spread about the world, including some people who are actually at the site of the news event itself.

4Additionally, there are web sites which fully document the details of these legends, and attempt to dissuade their dissemination, when harmful, as in the case of Craig Shergold

Jokes had not changed in many substantial ways with the growth of technology, other than adapting their transmission to new oral forms, until the advent of the Internet.

With the Internet, an essentially new form of oral joke arose: the “funny post”. The one particular form discussed here are the “True” story posts. These range from the “Darwin Awards”, an award given to a person who did the most amazingly stupid, funny and fatal thing in a given year, the “True Church Bulletin Quotes”, which contains a variety of true church bulletin errors and amusing contrasts, and the “True Foreign Signs” post, which contains amusingly mangled English.

Conclusion

Culturally owned narratives are still rather ubiquitous even in our mass media society. These narratives are often, however, not recognized as such, but would not likely otherwise be spread. Computer mediated communication is bringing new oral media into popular use, in email, faxes, chat groups, and other discussion forums. Other oral media, however, are still widely and effectively used in society to distribute culturally owned narratives.